# Entangled photon pair excitation and time-frequency filtered multidimensional photon correlation spectroscopy as a probe for dissipative exciton kinetics

### v1

@@ III.3 Multipoint photon correlation spectroscopy
In Fig. 7, we present six two-dimensional plots illustrating the signal where the filtering widths and centering parameters (the latter relates to the transport times) are varied. The centering times of the spectral filters are taken as the scanning parameters. In these plots, a series of features can be noticed: the peaks along the $\bar{\omega}_{1}$ axis correspond to transitions from the two to and one-exciton manifold ($\sim\omega_{fe}$), while the peaks along the $\bar{\omega}_{2}$ axis correspond to transitions from the one-exciton manifold to the ground state ($\sim\omega_{eg}$). The one-exciton manifold has a spectral width of $\approx 950\,\text{cm}^{-1}$. The density of energy gaps is significantly higher in the upper set of transitions. Even though, for the chosen example, the initial state distribution, following excitation of $f_{07}$, effectively spans only the first ten levels of the two-exciton manifold (an energy range of $\approx 350\,\text{cm}^{-1}$), the number of possible $\omega_{fe}$ transitions remains high.
@@
The upper-left panel (the reference simulation for this set) employs the following filtering parameters: spectral widths $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, temporal widths $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, and two-exciton (one-exciton) waiting times $t_{\text{w},1}=0\,\text{fs}$ ($t_{\text{w},2}=100\,\text{fs}$). Variables $t_{\text{w},j}$ are associated with the $t_{2},t_{1}$ integrals when the detector spectrograms are inserted into the relevant expression. We observe eight resonance zones spread along the $\bar{\omega}_{1}(/10^{3}\,\text{cm}^{-1})$ axis, each containing several resonances that can be resolved along the $\bar{\omega}_{2}(/10^{3}\,\text{cm}^{-1})$ axis. For easier interpretation, we may classify these resonances along $\bar{\omega}_{1}$ as four distinct clusters, containing the following number of resonance zones: two ($\sim 14.0-14.2$), three ($\sim 14.3-14.7$), two ($\sim 14.9-15.1$), and one ($\sim 15.3-15.5$). Each of these clusters may have multiple origins in the upper manifold. The four observed $\bar{\omega}_{1}$ resonance clusters correspond approximately to the transitions $\omega_{f_{1}e_{13}}$ and $\omega_{f_{2}e_{10}}-\omega_{f_{2}e_{13}}$; $\omega_{f_{1}e_{8}},\omega_{f_{2}e_{6}},\omega_{f_{2}e_{5}}$; $\omega_{f_{4}e_{3}}$ and either $\omega_{f_{3}e_{1}}$ or $\omega_{f_{4}e_{1}}$; and $\omega_{f_{7}e_{1}}-\omega_{f_{8}e_{1}}$, respectively. Taking each cluster, the $\bar{\omega}_{2}$ ($10^{3}\,\text{cm}^{-1}$) resonances are located at $\sim 15.5,15.8$ and $\sim 15.4,15.7$ (corresponding to $\omega_{e_{8}g},\omega_{e_{12}g}$ and $\omega_{e_{8}g},\omega_{e_{12}g}$); $\sim 15.3,15.6$, $\sim 15.3,15.4$, and $\sim 15.3,15.4$ ($\omega_{e_{7}g},\omega_{e_{8}g}$; $\omega_{e_{7}g},\omega_{e_{8}g}$; and $\omega_{e_{7}g},\omega_{e_{8}g}$); $\sim 15.1,15.3$ ($\omega_{e_{1}g},\omega_{e_{3}g}$); and $\sim 15.0$ ($\omega_{e_{1}g}$), respectively. This plot will serve as a reference simulation.
@@
The upper-right panel presents the results of varying the spectral filter width by setting $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$, while keeping all other parameters fixed to those of the reference simulation. We observe that several resonance zones in the second cluster are absent along the $\bar{\omega}_{1}$ axis, with the most noticeable absences occurring in the middle and lower energy sectors. Additionally, among the present zones, a few resonances along the $\bar{\omega}_{2}$ axis are absent, and the fourth cluster has been diminished in intensity. This observation may be rationalized by noting the increased destructive interferences permitted by the broadband spectral filter.
The middle-left panel investigates the effect of longer waiting times, specifically using $t_{\text{w},2}=1000\,\text{fs}$ for the one-exciton transport, while keeping all other parameters unchanged from the reference simulation. The population distribution in the one-exciton manifold is dependent on the cascaded emission from the two-exciton manifold. We note that the clusters have further coalesced along the $\bar{\omega}_{1}$ axis and are clearly separated into two dominant regions. The latter is indicative of the dipole transition rule dominating the spectral weight at these longer times. Furthermore, the fourth cluster has vanished, and the resolution along the $\bar{\omega}_{2}$ axis is lost. The middle-right panel explores the variation of resolution compared to the middle-left one by varying both the spectral filter width and the one-exciton waiting time ($t_{\text{w},2}$). The parameters are set jointly equal to those used in the two previous simulations ($\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ and $t_{\text{w},2}=1000\,\text{fs}$), while keeping all other parameters fixed. We observe that widening the spectral gate may not have permitted a larger number of resonance structures along the $\bar{\omega}_{2}$ axis. This implies that a broader set of transitions remain unaccounted for (either not presented or not dipole allowed). Importantly, these transitions, despite differing widely in energy, participate in similar dynamics due to the longer transport window afforded by the extended waiting time.
The lower panels are dedicated to exploring two distinct cases: the bottom-left panel examines the role of the width of the temporal filter (by setting $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$ ), and the bottom-right panel examines the role of the two-exciton transport waiting time (by setting $t_{\text{w},1}=50\,\text{fs}$), while keeping other parameters the same as the reference. For the bottom-left panel, we note that the second cluster has a diminished presence, with several resonances along the $\bar{\omega}_{2}$ axis being absent, an effect attributed to the shorter temporal filter effectively skipping some transitions occurring at longer respective timescales. For the bottom-right panel, in contrast with the reference simulation, we observe that the double resonance structures along the $\bar{\omega}_{2}$ axis have coalesced into single structures, while there is no significant change along the $\bar{\omega}_{1}$ axis. This can be attributed to the fact that at longer two-exciton transport duration, state-specific transport rates exhibited minor differences. The redistributed the spectral weights, in turn, has resulted in a fewer $\omega_{fe}$ transitions becoming dominant.
The initial states in all cases can be tracked from Figure 6. However, the state indices require a conversion to the energy domain (energy gaps) for constructing a robust phenomenological description.
@@ IV Conclusions
In this article, we had demonstrated that a combined protocol utilizing entangled photon-pair excitation and temporally-spectrally filtered photon counting provides a systematic route to probe dissipative multi-exciton kinetics. The protocol, illustrated via simulations for the exciton aggregate LHCII, is presented across three subsequent sections. It revealed, in section III.1, that narrow-band distribution around a special class of two exciton states can indeed be excited with considerable degree of control. The final population distribution is benchmarked against the variation of the entangled photon parameters in section III.1.2. Here we examined the pathway selectivity by targeting two different two exciton state that are excited via two pre-selected one-exciton states. The exciton-phonon interactions in multi-site molecular aggregates are characterized by exciton transport and dephasing dynamics spanning multiple time and energy scales. The entangled photon pairs demonstrated a superior ability maintain pathway selectivity. Following the excitation, and population evolution, illustrated in section III.2.1-III.2.2), we introduced the detection scheme to monitor the fate of the time-evolving population, and classify the emission pathways. It was shown, in section III.3, that the joint spectral and temporal filtering of the emitted fluorescence allow a more nuanced monitoring of inter-manifold transitions competing with intra-manifold transport kinetics.
We note that the targeted states, in III.1.2, are chosen due to their higher overlap with the local overtone states. These properties can be quantified by analyzing the transformation matrix elements, $T^{(2)}_{k,mm}$. While state-selective excitations in interacting molecular aggregates conventionally yield spatially delocalized wavepackets, an important corollary suggests that a specially crafted exciton wavepacket, that is, a tailored superposition of exciton eigenstates, may allow for the excitation of spatially localized wavepackets. Such localization could then enhance the propensity of exciton-exciton annihilation provided that significant non-adiabatic coupling elements between the higher-lying states allow the internal conversion to occur.
The expression in Eq. III.1 remains valid for a wide range of correlated optical fields, albeit in the weak-field regime. In the expression, five terms appear, each of which contains a convolution of inter-manifold coherence and intra-manifold population Green’s functions with four-point photon correlation functions. The latter can be interpreted as pathway weights, and multiobjective optimization schemes can be used to actively design them. Future work will address this issue.
Similarly, the expressions in Eq. III.3 remain valid for any general initial state distributions. The variation of the filtering parameters offers us the ability to classify inter-manifold transitions in the presence and competing with intra-manifold transport kinetics. It allows better assignment timescales and generate a dynamical picture for one and two-exciton transport kinetics. We assumed that the excitation and emission events are temporally separated, omitting spontaneous Raman contributions.
From these simulations, a probing strategy may emerge: preferentially accessing the one-exciton resonances that are less susceptible to transport within a given temporal window may be the key to achieving controlled excitation. We note that the full protocol encountered transport events thrice during its course. Within the model and the chosen expressions, the entanglement time, controlled by the length of the birefringent crystal, determines the temporal window of the photon-exciton interactions. The bandwidth of the classical SPDC pump, on the other hand, determines the range of accessible resonances. The independent controllability of these variables affords some flexibility in selecting their values. In turn, it will allow control over a broad range of time and frequency parameter domain. Analogously, the choice of filtering parameters, decided independently, must take into account the excitation properties to be able to characterize a wide range of two-photon emission pathways.
The utility of correlated photonic sources for multiphoton spectroscopies promises two main advantages in the appropriate parameter regime: the ability to actively modulate multiple material pathways to highlight unseen spectral signatures, and favorable scaling of the signal with the incoming field intensity. However, these benefits face a significant challenge due to the inherently weak signal-to-noise ratio. Technological advances in this direction would make the proposed protocol a suitable investigative tool for exploring potential non-classical effects in light-harvesting complex. Alternatively, the scheme remains applicable in the high-gain regime using squeezed photon sources.
@@ Appendix A The Hamiltonian
Here, $\lambda_{0}$ and $\gamma_{0}$ represent the damping strength and relaxation parameter for the overdamped mode, respectively. The Brownian oscillators are characterized by the Huang-Rhys parameter, mode frequency, and correlation time, denoted as $\lambda_{j}$, $\gamma_{j}$, $\omega_{j}$, respectively. Relevant parameter values are detailed in the [9, 8].
@@
The phonon interaction gives rise to exciton transport (first term) and dephasing (second term) Eq. II. The exciton transport components can be expressed as
$[\exp{(-K^{(n)}t)}]_{a_{1}a_{1},a_{3}a_{3}}$ $=\sum_{p}\chi_{a_{1}p}^{R}D_{pp}^{-1}e^{-\lambda_{p}t}\chi_{pa_{3}}^{L}$ (21)
@@
where $\lambda_{p}$ is the p-th eigenvalue of the transport matrix, $\chi^{L}(\chi^{R})$ are the left (right) eigenvectors $K\chi^{R}=\lambda_{p}\chi^{R}$ and $\chi^{L}K=\lambda_{p}\chi^{L}$ of the same with additional relation $D_{pp}=\chi^{L}\chi^{R}$. The inter-manifold coherence green’s functions, contains excitonic resonances $i\omega_{a_{1}a_{2}}^{(n)}-\gamma_{a_{1}a_{2}}^{(n)}t=z_{a_{1}a_{2}}^{(n)}$, where the line-broadening function term is evaluated as
$\gamma_{a_{j}}^{(n)}(t)$ $=\sum_{a_{j^{\prime}}}C_{\mathrm{ph}}(\omega_{a_{j}a_{j^{\prime}}}^{(n)})\sum_{m}T_{ma_{j}}^{(n)}T_{ma_{j}}^{(n)}T_{ma_{j^{\prime}}}^{(n)}T_{ma_{j^{\prime}}}^{(n)}$ (22)
@@
and the phonon correlation function evaluated at resonant exciton transition frequencies is given by
$C_{\text{phonon}}(\Omega)=\int_{0}^{\infty}dt\exp(i\Omega t)$
$\qquad\times\int\frac{d\omega}{2\pi}J(\omega)\Big[\coth{(\beta\omega/2)}\cos{\omega t}\mp i\sin{\omega t}\Big]$ (23)
Therefore characteristic broadening of the resonances is determined by the phonon spectral function, and exciton gaps.
Appendix B Note on the observable: two-exciton population distribution

### v2

@@ III.3 Photon correlation spectroscopy
In Fig. 8, we present six two-dimensional plots where centering times of the frequency filters (i.e., $\bar{\omega}_{j}$) are taken as the scanning parameters. These results explore the signal for cases where the widths (i.e., $\sigma_{\bar{t}_{j}}$) and centering parameters of time filters (i.e., $\bar{t}_{j}$) are varied. We note that, the latter help unravel the exciton transport timescales. The two-exciton distribution targeted at $f_{07}$, after a period of time evolution serves as the initial configuration. The peaks along the $\bar{\omega}_{1}$ axis correspond to transitions from the two-exciton to and one-exciton manifold $|f\rangle\rightarrow|e\rangle$, while the peaks along the $\bar{\omega}_{2}$ axis correspond to transitions from the one-exciton manifold to the ground state $|e\rangle\rightarrow|g\rangle$. The former monitors the spectral gaps $\sim\omega_{fe}$ while the latter the spectral gaps $\sim\omega_{eg}$. This can be ensured by choosing the centering times of the time filter $\bar{t}_{2}$ well-ahead of $\bar{t}_{1}$.
As highlighted earlier in Section II, density of energy gaps is significantly higher in the upper set of transitions. We recall that the one-exciton manifold hosts $14$ states within a spectral width of $\approx 970\,\text{cm}^{-1}$ and following excitation of $f_{07}$, the two-exciton manifold hosts $\sim 10$ states within a spectral width of $\approx 350\,\text{cm}^{-1}$. The number of plausible transitions remains large $\sim 140$. Two filtering stages allow classify these transitions in different temporal windows.
@@
For the upper-left panel, the reference simulation for this set, we use the following parameters: $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, waiting times $t_{\text{w},1}=0\,\text{fs}$, and $t_{\text{w},2}=100\,\text{fs}$. Variables $t_{\text{w},j}$ are associated with the $\tau_{j}$ integrals in the expression of detector spectrograms. The waiting times determine the duration for which exciton transport is allowed at each manifold, before the time-filter is opened. In the plot, we observe eight resonance zones spread along the $\bar{\omega}_{1}$ axis, each containing several resonances that can be further resolved along the $\bar{\omega}_{2}$ axis. For easier interpretation, we may classify the resonances along $\bar{\omega}_{1}/10^{3}\,\text{cm}^{-1}$ in four distinct clusters, each containing two ($\sim 14.0-14.2$), three ($\sim 14.3-14.7$), two ($\sim 14.9-15.1$), and one ($\sim 15.3-15.5$) resonances. Since only the transition energy differences are being captured, each of these transitions, in the absence of the time filter would have appeared in a congested form. The four observed clusters correspond approximately to the transitions $\omega_{f_{1}e_{13}}$ and $\omega_{f_{2}e_{10}}-\omega_{f_{2}e_{13}}$; $\omega_{f_{1}e_{8}},\omega_{f_{2}e_{6}},\omega_{f_{2}e_{5}}$; $\omega_{f_{4}e_{3}}$ and either $\omega_{f_{3}e_{1}}$ or $\omega_{f_{4}e_{1}}$; and $\omega_{f_{7}e_{1}}-\omega_{f_{8}e_{1}}$, respectively. For each of these clusters, the resonances along $\bar{\omega}_{2}/10^{3}\,\text{cm}^{-1}$ are identified as $\sim 15.5,15.8$ and $\sim 15.4,15.7$ (corresponding to $\omega_{e_{8}g},\omega_{e_{12}g}$ and $\omega_{e_{8}g},\omega_{e_{12}g}$); $\sim 15.3,15.6$, $\sim 15.3,15.4$, and $\sim 15.3,15.4$ ($\omega_{e_{7}g},\omega_{e_{8}g}$; $\omega_{e_{7}g},\omega_{e_{8}g}$; and $\omega_{e_{7}g},\omega_{e_{8}g}$); $\sim 15.1,15.3$ ($\omega_{e_{1}g},\omega_{e_{3}g}$); and $\sim 15.0$ ($\omega_{e_{1}g}$), respectively.
@@
The upper-right panel repeats the simulation with $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$, while keeping all other parameters fixed to those of the previous one. We observe, along the $\bar{\omega}_{1}$ axis, that several resonance zones in the second cluster are absent; the most noticeable absences occur in the middle and low energy sectors. Additionally, a few resonances along the $\bar{\omega}_{2}$ are absent. The fourth cluster has been diminished in intensity. The broadband spectral filters used in the simulation increases the number of destructive interferences. Given that the transitions originate from a limited number of states number of states it also highlights that they are preferentially transitioning to a few selected one-exciton states.
The middle-left panel investigates the effect of longer waiting times aimed at identifying the role of one-exciton transport. We set $t_{\text{w},2}=1000\,\text{fs}$ while keeping all other parameters unchanged from the reference simulation. The population distribution in the one-exciton manifold depends on two factors: the dipole-connectivity of states with those of the two-exciton manifold and the transport rates. We note that along the $\bar{\omega}_{1}$ axis the clusters have further coalesced. They are clearly separated into two dominant regions. Furthermore, the fourth cluster has vanished, and the resolution along the $\bar{\omega}_{2}$ axis is lost. The observation indicates that, at longer waiting times, the connectivity dictated by transition dipoles focuses the spectral weights to a few selected states.
The middle-right panel repeats the reference simulation but with $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ and $t_{\text{w},2}=1000\,\text{fs}$. Therefore it is aimed at exploring the role of a broadband frequency filter on one-exciton transport at longer times. We note the resonances along the $\bar{\omega}_{2}$ axis and conclude that widening the spectral gate may not have permitted a larger number of transitions. The reasons are twofold: the initial two-exciton distributions, being restricted to a limited set of states, have no further transitions to explore and these transitions at longer times participates in dynamics that converges to same set of states.
The bottom-left panel explores the examines the role of the width of the time filter by setting $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$ while keeping all other parameters the same as the reference. We note that the second cluster has a diminished presence. Also, several resonances along the $\bar{\omega}_{2}$ axis are absent. We note that positioning the centering time of the filter at shorter time effectively skips the resonances that are generated as a result of the longer time dynamics. Hence, by comparing the reference simulation and this one, one may identify the transitions that takes place at longer times.
The bottom-right panel examines the role of the two-exciton transport by setting $t_{\text{w},1}=50\,\text{fs}$. All other parameters are kept as the same as that of the reference. We observe that the double resonance structures along the $\bar{\omega}_{2}$ axis have coalesced into single structure, while there is no significant change along the $\bar{\omega}_{1}$ axis. The longer two-exciton transport tend to funnel all populations to a few low-lying two-exciton states, as can be observed in Fig. 5 and Fig. 7. This makes a fewer number of transitions possible, reducing the number of corresponding resonances. We note that the for a robust phenomenological description the state indices from the figures need to be converted to energy values.
@@ IV Conclusions
In this work, we demonstrated that a combining entangled photon-pair excitation and time-frequency filtered two-photon counting provides a systematic route to probe dissipative one- and two-exciton kinetics. The combined protocol was applied to the interacting molecular aggregate, light-harvesting complex II (LHCII), where exciton–exciton and exciton–phonon interactions jointly govern to the dynamics. The protocol offers a classification scheme for exciton dynamics, even when they span multiple time and frequency scales.
Results in the Section III.1 reveals that a narrow-band population distribution around some of the two-exciton states can be created with a high degree of fidelity. The final population distribution is examined against variations in the entangled photon parameters and the temporal width of the SPDC pump. We also examined the differential nature of excitation by targeting two distinct two-exciton states. Because these two target states are excited via the same one-exciton states, they offer insight into pathway selectivity. From the current simulations, a strategy emerges: preferentially accessing the one-exciton resonances that are less susceptible to transport is the key to achieving robust, state-selective excitation of two-exciton states. The expression in Eq. III.1 remains valid for a wide range of photonic sources, provided the dynamics remains within the perturbative regime of exciton–photon coupling. The expression consists of five terms, each containing a convolution of exciton Green’s functions with four-point correlation functions of the driving photonic sources. They modulate spectral weights for each pathway. Our future work will address how multiobjective optimization algorithms can be used to actively tailor these correlation functions.
The evolution of the initially created two-exciton population is studied in Section III.2. This evolved population distribution serves as the initial configuration for the photon emission pathways. The presence of transport within the two-exciton manifold effectively scrambles the identity of the states initially created by two-photon absorption. Similarly, transport in the one-exciton manifold have the same effects on the state distribution generated after the first photon emission. Interestingly, the one-exciton transport may exert the scrambling effect on two occasions: during the excitation and again during the emission process.
In Section III.3, we deployed the time-frequency filtering to track inter-manifold transitions. The expressions in Eq. III.3 remain valid for general two-exciton population distributions. The variation of the filtering parameters may offer detailed insights into spectral gaps, intra-manifold transport timescales, and inter-manifold dephasing. We assumed that the excitation and emission events are temporally separated which omits spontaneous Raman contributions. The joint spectral-temporal filtering effectively enhances the resolution of the bare two-photon counting signal. Two spectrograms, parametrized by filtering functions separate the exciton pathways jointly in time and frequency. It allows one to construct a kinetic description of state-resolved exciton dynamics even in the presence of dissipation.
The entanglement time, controlled by the length of the birefringent crystal, determines the temporal window of successive photon-exciton interactions. In contrast, the bandwidth of the classical spontaneous parametric down-conversion (SPDC) pump determines the range of accessible resonances. The independent controllability of these variables may affords additional flexibility. For a given $N$-site exciton aggregate, there are $N$ one-exciton states and N+N(N1)/2N+N(N-1)/2 two-exciton states. As the size of the aggregate increases, the dynamics may span a broad range of time and frequency values. To probe such dynamics a wider tunability of photonic parameters may become necessary [5, 37, 11]. Similar constraints apply on the filtering parameters, which must be able to characterize a wide range of two-photon emission pathways.
In general, entangled photon pairs offer two primary advantages for two-photon spectroscopy: the ability to modulate competing excitation (and de-excitation) pathways to selectively suppress or amplify specific exciton resonances, and favorable scaling of the signal with the intensity of the incoming field. We note that the correlation function of coherent laser fields factorizes into a product of four amplitude functions, i.e., $\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle=A^{*}_{4}(\omega_{4})A^{*}_{3}(\omega_{3})A_{2}(\omega_{2})A_{1}(\omega_{1})$. The signal scales with the square of the incident field intensity. In contrast, for the entangled photon pairs, the signal scales linearly. However, realizing these benefits remains a significant challenge due to the inherently low signal-to-noise ratio and the difficulty of separating contributions arising from photonic entanglement [45, 26, 27]. Technological advances in this direction would establish the proposed protocol as a suitable tool for investigating non-classical effects in light-harvesting complexes.
The time- and frequency-filtered $n$-photon counting signal can analyze multiphoton emission in a $2n$-dimensional parameter space. If combined with an efficient state preparation using multiphoton entangled sources, it may provide a robust spectroscopic tool for characterizing dissipative molecular excitonic systems. Such approaches can also be extended to programmable photonic sources and interferometric detection schemes [28, 13, 3].
@@ Appendix A The Hamiltonian
Here, $\lambda_{0}$ and $\gamma_{0}$ represent the damping strength and relaxation parameter for the overdamped mode, respectively. The Brownian oscillators are characterized by the Huang-Rhys parameter, mode frequency, and correlation time, denoted as $\lambda_{j}$, $\gamma_{j}$, $\omega_{j}$, respectively. Relevant parameter values are detailed in the [10, 9, 41, 40].
@@
The phonon interaction gives rise to exciton transport (first term) and dephasing (second term) Eq. II. The exciton transport Green’s function can be expressed as
$\mathcal{G}_{a_{1}a_{1},a_{3}a_{3}}(t)$ $=[\exp{(-Kt)}]_{a_{1}a_{1},a_{3}a_{3}}$
$=\sum_{p}\chi_{a_{1}p}^{R}D_{pp}^{-1}e^{-\lambda_{p}t}\chi_{pa_{3}}^{L}$ (19)
@@
where $\lambda_{p}$ is the p-th eigenvalue of the transport matrix, $\chi^{L}(\chi^{R})$ are the left (right) eigenvectors that satisfy $K\chi^{R}=\lambda_{p}\chi^{R}$ and $\chi^{L}K=\lambda_{p}\chi^{L}$ of the same with additional relation $D=\chi^{L}\chi^{R}$.
The intra-manifold exciton coherence Green’s function is given by
$\mathcal{G}_{a_{1}a_{2}}(t)$ $=\exp{[-i\omega_{a_{1}a_{2}}t-\gamma_{a_{1}a_{2}}t]}$ (20)
where we defined the pure dephasing parameter
$\tilde{\gamma}_{a_{1}a_{2}}=\gamma_{a_{1}a_{2}}-\frac{1}{2}(K_{a_{1}a_{1},a_{1}a_{1}}+K_{a_{2}a_{2},a_{2}a_{2}}).$ (21)
and the dephasing parameters $\gamma_{a_{1}a_{2}}=-K_{a_{1}a_{2},a_{1}a_{2}}$.
Figure 9: Histograms describing distribution of spectral gaps corresponding for $\omega_{eg}$, $\omega_{fg}$, and $\omega_{fe}$ at the top, middle and bottom panel respectively. Each energy bin have a width of $200\,\text{cm}^{-1}$. An increase in the number of counts indicates corresponding increase in inter-manifold transitions. For a detailed discussion, refer to Appendix B.
The state-specific line-broadening functions are defined as
$\gamma_{a_{1}}(t)$ $=\sum_{a_{2}}C(\omega_{a_{1}a_{2}})\sum_{m}T_{ma_{1}}T_{ma_{1}}T_{ma_{2}}T_{ma_{2}}$ (22)
@@
The dissipation function, $C(\omega_{ab})$, is the Fourier transform of the phonon correlation function evaluated at exciton transition frequencies $\omega_{ab}$, expressed as
$C(\omega_{ab})=\int_{0}^{\infty}dt\exp(i\omega_{ab}t)$
$\qquad\times\int\frac{d\omega}{2\pi}J_{\text{ph}}(\omega)\Big[\coth{(\beta\omega/2)}\cos{\omega t}\mp i\sin{\omega t}\Big]$ (23)
It indicates that the line broadening functions are determined by the phonon spectral function, and exciton gaps. The inter-manifold coherence Green’s functions, contains excitonic resonances of the type: $iz_{a_{1}a_{2^{\prime}}}=\omega_{a_{1}a_{2^{\prime}}}-i\gamma_{a_{1}a_{2^{\prime}}}t$, where the primed indices indicate states belonging to different manifolds. The values of the inter-manifold dephasing are estimated from the expression of the line broadening by replacing respective manifold terms with primed indices. Correlations between the source of driving and exciton-phonon dissipation was neglected [12].
Appendix B Note on the exciton energy distribution: one and two-exciton manifold
As stated in section II), the exact-diagonalization step yields $14$, one exciton states (denoted $|e_{j}\rangle$) and $105$ two-states (denoted $|f_{k}\rangle$). The distribution of transition energy gaps or spectral gaps for each of these manifolds can be evaluated using a pair of exciton energies belonging to different manifolds. For transitions between the ground and one-exciton states, ranging from $\omega_{e_{1}g_{1}}=E_{e_{1}}$ to $\omega_{e_{14}g_{1}}=E_{e_{14}}$, there are $14$ spectral gaps. The same between the ground and two-exciton states are $105$, which ranges from $\omega_{f_{1}g_{1}}=E_{f_{1}}$ to $\omega_{f_{105}g_{1}}=E_{f_{14}}$. In comparison, spectral gaps between the one and two exciton manifold spans from $\omega_{f_{1}e_{14}}=13909.6\,\text{cm}^{-1}$ to $\omega_{f_{105}e_{1}}=16847.3\,\text{cm}^{-1}$ contributing to a $14\times 105$, a polynomial large number of spectral gaps. A majority of them are dipole-allowed. These gaps appear in the expressions of exciton resonances via Green’s functions. The distribution of spectral gaps are visualized via three histograms in Fig. 9.
In the Fig. 4, the simulation considered a special situation: the pump photon energies are downconverted to degenerate entangled photon pairs. The energies of the sequentially corresponds to the each spectral gaps corresponding to each state in the the two-exciton manifold. For the chosen two examples of two exciton states $f$ and $f$, described in
Appendix C Note on the observable: two-exciton population distribution